boat moves upstream in 6 hrs and covers the same distance downstream in 5 hrs. then speed of a raft floating?

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

5=d/v+vr

6=d/v-vr

divide both

5/6= v-vr/v+vr

solving

vr=v/11

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..